# On Mirzakhani’s twist torus conjecture

### v2

@@ 4.2. Affine invariant subvarieties
Invoking Theorem 4.2, this implies that the Fölner sets $\{g_{t}u_{s}q:t\in[0,T],s\in[0,r]\}$ based at any $q\in\mathcal{Q}\mathbb{T}^{0}$ equidistribute to $\nu_{\mathcal{L}}$ as $T\to\infty$. Integrating against the twist torus measure and applying Lemma 4.1 yields
[…]
for every $f\in C_{c}(\mathcal{Q})$, where $\mathcal{Q}$ is the ambient stratum. This is what we wanted to show.
In general, the twist torus $\mathcal{Q}\mathbb{T}^{0}$ may not be minimal for the horocycle flow. In this case, we show that the $P$–orbit closure of $\nu^{0}$–almost every point is $\mathcal{L}$; the rest of the argument then proceeds as above.
On any $q\in\mathcal{Q}\mathbb{T}^{0}$, choose a basis for relative homology consisting of horizontal saddle connections and saddles which connect the boundary components of the cylinders of $q$; this gives a local period coordinate chart for the stratum $\mathcal{Q}$. Fix a finite collection121212 In fact, one can cover a lift of the twist torus to $\mathcal{Q}^{1}\mathcal{T}_{g}$ with a single chart of this form, but we use a finite collection of restricted charts so that we do not need to worry about the $\operatorname{Stab}(\gamma)$ action. of such charts that covers $\mathcal{Q}\mathbb{T}^{0}$. Let $\mathcal{N}\subsetneq\mathcal{L}$ be any affine invariant subvariety not containing $\mathcal{Q}\mathbb{T}^{0}$. In any of the period charts constructed above, $\mathcal{Q}\mathbb{T}^{0}$ is cut out by affine equations of the form $\operatorname{hol}(s)=\ell_{i}\in\mathbb{R}$ (specifying the length of the horizontal saddle connections) and equations of the form $\operatorname{Im}(\operatorname{hol}(s))=h_{i}$ (specifying the heights of the cylinders); compare §4.1. On the other hand, in period coordinates $\mathcal{N}$ is a union of finitely many linear subspaces, each of which is cut out by linear equations with real coefficients. Therefore, we see that in each chart, $\mathcal{Q}\mathbb{T}^{0}\cap\mathcal{N}$ is a union of finitely many proper subspaces, hence has $\nu^{0}$ measure 00. Moreover, we note that each subspace must be tangent to a proper subtorus of $\mathcal{Q}\mathbb{T}^{0}$, else its closure would be the entirety of $\mathcal{Q}\mathbb{T}^{0}$.
As the $P$–orbit closure of any $q$ is always an AIS and there are only countably many such (Theorem 4.2.1), we see that
$\{q\in\mathcal{Q}\mathbb{T}^{0}:\overline{Pq}\neq\mathcal{L}\}$
is a union of countably many proper subtori of $\mathcal{Q}\mathbb{T}^{0}$. In particular, this set has measure zero with respect to $\nu^{0}$, and so the $P$–orbit closure of almost every point of $\mathcal{Q}\mathbb{T}^{0}$ is $\mathcal{L}$. ∎
@@ 4.3. Individual pushforwards
Let $\mathcal{Q}$ be a stratum of quadratic differentials and suppose that $\nu$ is a $U$-invariant probability measure on $\mathcal{Q}$. If the limit
[…]
@@
$(g_{t})_{*}\nu_{\gamma}(\mathsf{Y},\mathbf{h})\to\nu_{\mathcal{L}}$
@@
Forni also proved that if $\nu$ is especially nice, then the set $Z$ is empty. Recall that a horospherical measure $\nu$ is one that is supported in a leaf of the unstable foliation of $g_{t}$ on a stratum $\mathcal{Q}\subset\mathcal{Q}^{1}\mathcal{M}_{g}$, is absolutely continuous with respect to the canonical affine measure on the leaf, has continuous density, and has conditional measures along horocycle orbits equal to the restriction of one-dimensional Lebesgue measures.131313 Theorem 1.6 in Forni’s paper is stated for the stable foliation, but clearly applies to the unstable.
@@
Theorem 1.6 of [For21] (or [EMM22, Proposition 3.2]) ensures $(g_{t}\circ\mathcal{O})_{*}\sigma$ converges weak-* to the Masur–Smillie–Veech measure $\nu_{\infty}$ on the ambient stratum as $t\to\infty$. Applying Theorem 1.4 then gives that $(\operatorname{Dil}_{t})_{*}\sigma$ converges weak-* to $\mathcal{O}^{*}\nu_{\mathcal{Q}}$, where $\mathcal{Q}$ is Masur–Smillie–Veech measure on the stratum component of $\mathcal{Q}^{1}\mathcal{M}_{g}$ containing the horosphere (slice). ∎
5. Identifying orbit closures
Our goal is now to identify $\mathcal{L}=\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\underline{A},\mathbf{h})}$. For this, we will need the notion of the rank of an AIS. Let $\widehat{g}\geq 2$ and suppose ^$\widehat{}\underline{\kappa}$ is a partition of $22\widehat{g}-$ 2. We consider the unit area locus $\widehat{\mathcal{L}}$ of an AIS in the stratum (^$\mathcal{H}(\widehat{}\underline{\kappa})$ of abelian differentials determined by ^$\widehat{}\underline{\kappa}$. Then $\mathbb{R}\widehat{\mathcal{L}}$ is an immersed submanifold of (^$\mathcal{H}(\widehat{}\underline{\kappa})$ that is locally a $\mathbb{C}$-linear subspace in period coordinates for (^$\mathcal{H}(\widehat{}\underline{\kappa})$ defined over $\mathbb{R}$ (see §4.2). For $\omega\in\mathbb{R}\widehat{\mathcal{L}}$, we have
$T_{\omega}\mathbb{R}\widehat{\mathcal{L}}\subset T_{\omega}\mathcal{H}(\widehat{}\underline{\kappa})\cong H^{1}(\omega,Z(\omega);\mathbb{C}),$
and there is a natural linear projection $p$ from relative to absolute cohomology.
@@ 5. Identifying orbit closures
The rank of $\widehat{\mathcal{L}}$ is the half the complex dimension of $p(T_{\omega}\mathbb{R}\widehat{\mathcal{L}})\subset H^{1}(\omega;\mathbb{C})$. An AIS is said to have high rank if its rank is at least $\widehat{g}/2+1$.
Recent deep work of Apisa and Wright classifies the AISs of high rank.

### v3

@@ 4.2. Affine invariant subvarieties
Invoking Theorem 4.4, this implies that the Fölner sets $\{g_{t}u_{s}q:t\in[0,T],s\in[0,r]\}$ based at any $q\in\mathcal{Q}\mathbb{T}^{0}$ equidistribute to $\nu_{\mathcal{L}}^{1}$ as $T\to\infty$. Integrating against the twist torus measure and applying Lemma 4.3 yields
$\frac{1}{T}\int_{0}^{T}\int_{0}^{r}\int_{\mathcal{Q}\mathbb{T}^{0}}f(g_{t}u_{s}q)\,d\nu^{0}(q)\,ds\,dt=\frac{1}{T}\int_{0}^{T}\int_{\mathcal{Q}\mathbb{T}^{t}}f(q)\,d\nu^{t}(q)\,dt\to\int_{\mathcal{L}^{1}}f\,d\nu_{\mathcal{L}}^{1}$
for every $f\in C_{c}(\mathcal{Q}^{1})$, where $\mathcal{Q}^{1}$ is unit area locus of the ambient stratum. This is what we wanted to show.
In general, the twist torus $\mathcal{Q}\mathbb{T}^{0}$ may not be minimal for the horocycle flow. In this case, we show that the PP–orbit closure of $\nu^{0}$–almost every point is $\mathcal{L}^{1}$; the rest of the argument then proceeds as above.
On any $q\in\mathcal{Q}\mathbb{T}^{0}$, choose a basis for relative homology consisting of horizontal saddle connections and saddles which connect the boundary components of the cylinders of $q$; this gives a local period coordinate chart for the entire stratum $\mathcal{Q}$ (not just its unit area locus). Fix a finite collection [^13] In fact, one can cover a lift of the twist torus to $\mathcal{QT}_{g}$ with a single chart of this form, but we use a finite collection of restricted charts so that we do not need to worry about the $\operatorname{Stab}(\gamma)$ action. of such charts that covers $\mathcal{Q}\mathbb{T}^{0}$. Let $\mathcal{N}\subsetneq\mathcal{L}$ be any affine invariant subvariety (again, not just its unit area locus) not containing $\mathcal{Q}\mathbb{T}^{0}$. In any of the period charts constructed above, $\mathcal{Q}\mathbb{T}^{0}$ is cut out by affine equations of the form $\operatorname{hol}(s)=\ell_{i}\in\mathbb{R}$ (specifying the length of the horizontal saddle connections) and equations of the form $\operatorname{Im}(\operatorname{hol}(s))=h_{i}$ (specifying the heights of the cylinders); compare §4.1. On the other hand, in period coordinates $\mathcal{N}$ is a union of finitely many linear subspaces, each of which is cut out by linear equations with real coefficients. Therefore, we see that in each chart, $\mathcal{Q}\mathbb{T}^{0}\cap\mathcal{N}$ is a union of finitely many proper subspaces, hence has $\nu^{0}$ measure 0. Moreover, we note that each subspace must be tangent to a proper subtorus of $\mathcal{Q}\mathbb{T}^{0}$, else its closure would be the entirety of $\mathcal{Q}\mathbb{T}^{0}$.
As the PP–orbit closure of any $q$ is always an AIS and there are only countably many such (Theorem 4.4.1), we see that
$\{q\in\mathcal{Q}\mathbb{T}^{0}:\overline{Pq}\neq\mathcal{L}^{1}\}$
is a union of countably many proper subtori of $\mathcal{Q}\mathbb{T}^{0}$. In particular, this set has measure zero with respect to $\nu^{0}$, and so the PP–orbit closure of almost every point of $\mathcal{Q}\mathbb{T}^{0}$ is $\mathcal{L}^{1}$. ∎
@@ 4.3. Individual pushforwards
Let $\mathcal{Q}^{1}$ be a stratum of unit area quadratic differentials and suppose that $\nu$ is a $U$-invariant probability measure on $\mathcal{Q}^{1}$. If the limit
$\nu_{\infty}=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}(g_{t})_{*}\nu\,dt$
@@
$(g_{t})_{*}\nu_{\gamma}(\mathsf{Y},\mathbf{h})\to\nu_{\mathcal{L}}^{1}$
@@
Forni also proved that if $\nu$ is especially nice, then the set $Z$ is empty. Recall that a horospherical measure $\nu$ is one that is supported in a leaf of the unstable foliation of $g_{t}$ on a stratum $\mathcal{Q}^{1}\subset\mathcal{Q}^{1}\mathcal{M}_{g}$, is absolutely continuous with respect to the canonical affine measure on the leaf, has continuous density, and has conditional measures along horocycle orbits equal to the restriction of one-dimensional Lebesgue measures.
@@
Theorem 1.6 of [For21] [^14] Theorem 1.6 in Forni’s paper is stated for the stable foliation, but clearly applies to the unstable. (or [EMM22, Proposition 3.2]) ensures $(g_{t}\circ\mathcal{O})_{*}\sigma$ converges weak-* as $t\to\infty$ to the Masur–Smillie–Veech probability measure $\nu_{\infty}^{1}$ on the ambient stratum component of $\mathcal{Q}^{1}\mathcal{M}_{g}$ containing the horosphere (slice). Applying Theorem 1.3 then gives that $(\operatorname{Dil}_{t})_{*}\sigma$ converges weak-* to $\mathcal{O}^{*}\nu_{\mathcal{Q}}^{1}$. ∎
5. Twist tori built from pants decompositions
In this section, we identify the orbit closures of twist tori built out of pants decompositions, proving Theorem 1.2. The key tools are some properties of affine invariant subvarieties and Wright’s notion of (cylinder) rank from [Wri15], discussed below. Throughout this section and the next, except when we discuss equidistribution, we will always work with an entire stratum/AIS, not just its unit area locus.
5.1. Cylinders and rank
Let $\mathcal{N}$ be an AIS in a stratum of (squares of) abelian differentials $\mathcal{H}=\mathcal{Q}\mathcal{M}_{g}(\underline{\kappa};+1)$. Then $\mathcal{N}$ is an immersed submanifold of $\mathcal{H}$ that is locally a $\mathbb{C}$-linear subspace in period coordinates for $\mathcal{H}$ defined over $\mathbb{R}$ (see §4.2). For $\omega\in\mathcal{N}$, we have
$T_{\omega}\mathcal{N}\leq T_{\omega}\mathcal{H}\cong H^{1}(\omega,Z(\omega);\mathbb{C}).$
There is a natural linear projection
$p:H^{1}(\omega,Z(\omega);\mathbb{C})\to H^{1}(\omega;\mathbb{C}).$
@@ 5.1. Cylinders and rank
The rank of $\mathcal{N}$ is the half the (complex) dimension of $p(T_{\omega}\mathcal{N})\leq H^{1}(\omega;\mathbb{C})$. The rel of $\mathcal{N}$ is the (complex) dimension of $T_{\omega}\mathcal{N}\cap\ker p$.
Rank and rel are among the most important invariants of an AIS. The proof of the following uses only the definitions and basic properties.
Lemma 5.2.
Suppose $\mathcal{N}$ and $\mathcal{N}^{\prime}$ are AISs in a stratum of squares of abelian differentials. Assume that $\mathcal{N}\subset\mathcal{N}^{\prime}$ and
$\textrm{rank}(\mathcal{N})+\textrm{rel}(\mathcal{N})\geq\textrm{rank}(\mathcal{N}^{\prime})+\textrm{rel}(\mathcal{N}^{\prime}).$
Then $\mathcal{N}=\mathcal{N}^{\prime}$.
Proof.
Let $\omega\in\mathcal{N}$ be a smooth point. [^15] It is possible that $\mathcal{N}$ is entirely contained in the orbifold locus of $\mathcal{H}$. In this case, by “smooth” we mean a point whose local orbifold group is the smallest possible. See, e.g., [CSW20, §2.3]. In any case, our arguments only use the dimension of the relevant spaces, not their vector space structure. Then $T_{\omega}\mathcal{N}\leq T_{\omega}\mathcal{N}^{\prime}.$ Directly from this containment, we deduce
$\textrm{rank}(\mathcal{N})\leq\textrm{rank}(\mathcal{N}^{\prime})\text{ and }\textrm{rel}(\mathcal{N})\leq\textrm{rel}(\mathcal{N}^{\prime}).$
By hypothesis and the fact that rank and rel are never negative, we conclude
$\textrm{rank}(\mathcal{N})=\textrm{rank}(\mathcal{N}^{\prime})\text{ and }\textrm{rel}(\mathcal{N})=\textrm{rel}(\mathcal{N}^{\prime}).$
The (complex) dimension of $\mathcal{N}$ is computed as
$\dim(\mathcal{N})=2\textrm{rank}(\mathcal{N})+\textrm{rel}(\mathcal{N}),$
and similarly for $\mathcal{N}^{\prime}$. Thus $\dim(\mathcal{N})=\dim(\mathcal{N}^{\prime})$, which implies that $\mathcal{N}=\mathcal{N}^{\prime}$ near $\omega$. By $\mathsf{GL}(2,\mathbb{R})$-invariance, $\mathcal{N}=\mathcal{N}^{\prime}$ globally. ∎
Wright introduced the notion of rank in [Wri15] and related it to the dimension of the space spanned by the directions tangent to certain cylinder deformations, which we recall below.
For a given $\omega\in\mathcal{N}\subset\mathcal{H}$, let $\mathcal{C}(\omega)$ be the set of maximal, singularity free horizontal cylinders in $\omega$. Each element $C_{i}\in\mathcal{C}(\omega)$ is an isometrically embedded open Euclidean cylinder with positive height, foliated by closed, nonsingular oriented leaves of the imaginary foliation of $\omega$ which are all homotopic to its core curve $\gamma_{i}$. Recall that
$u_{s}=\begin{pmatrix}1&s\\ 0&1\end{pmatrix}\in\mathsf{SL}_{2}\mathbb{R},$
and define $u_{s}^{C_{i}}q\in\mathcal{H}(\underline{\kappa})$ to be the abelian differential obtained by applying the horocycle flow $u_{s}$ to $C_{i}$ but not the rest of $\omega$. Denote by $\eta_{i}\in T_{\omega}\mathcal{H}(\underline{\kappa})$ the tangent vector to the path $s\mapsto u_{s}^{C_{i}}\omega$ at $s=0$, and observe that $\eta_{i}\in H^{1}(\omega,Z(\omega);\mathbb{R})$. In fact, $\eta_{i}$ is Poincaré–Lefschetz dual to the homology class in $H_{1}(\omega\setminus Z(\omega);\mathbb{R})$ defined by $h_{i}\gamma_{i}$, where $h_{i}$ is the height of the cylinder $C_{i}$; see [MW18, §4].
The fact that $\mathcal{N}$ is cut out by $\mathbb{R}$-linear equations implies $T_{\omega}\mathcal{N}=T_{\omega}^{\mathbb{R}}\mathcal{N}\otimes\mathbb{C}$, where $T_{\omega}^{\mathbb{R}}\mathcal{N}\leq H^{1}(\omega,Z(\omega);\mathbb{R})$. The twist space of $\mathcal{N}$ at $\omega$ is spanned by those cylinder deformations which remain in $\mathcal{N}$:
$\mathrm{Twist}(\omega,\mathcal{N}):=\langle\eta_{i}:C_{i}\in\mathcal{C}(\omega)\rangle\cap T_{\omega}^{\mathbb{R}}\mathcal{N}.$
We deduce the following from [Wri15, Section 8].
Proposition 5.3.
Let $\mathcal{N}$ be an AIS in a stratum of squares of abelian differentials and suppose $\omega\in\mathcal{N}$ is horizontally periodic. Then
$\dim\mathrm{Twist}(\omega,\mathcal{N})\leq\textrm{rank}(\mathcal{N})+\textrm{rel}(\mathcal{N}),$
with equality if no other $\omega^{\prime}\in\mathcal{N}$ has more cylinders than $\omega$.
Proof.
Corollary 8.11 of [Wri15] asserts that, for any $\omega\in\mathcal{N}$,
$\dim T_{\omega}^{\mathbb{R}}\mathcal{N}-\dim\mathrm{Twist}(\omega,\mathcal{N})\geq\textrm{rank}(\mathcal{N}).$
Combining Lemma 8.6 and Lemma 8.8 of [Wri15] gives equality if $\omega$ has the maximal number of horizontal cylinders among all $\omega^{\prime}\in\mathcal{N}$. Using $\dim T_{\omega}^{\mathbb{R}}\mathcal{N}=2\textrm{rank}(\mathcal{N})+\textrm{rel}(\mathcal{N})$, we obtain the desired conclusion. ∎
Lemma 8.11 of [Wri15], used in the proof of Proposition 5.3, relies on the fact that $p(T_{\omega}^{\mathbb{R}}\mathcal{N})$ is a symplectic subspace of $H^{1}(\omega;\mathbb{R})$ with respect to the intersection pairing [AEM17, Theorem 1.4]. The rank can then be interpreted as the (real) dimension of a Lagrangian subspace of $p(T_{\omega}^{\mathbb{R}}\mathcal{N})$. Since $\eta_{i}$ is Poincaré–Lefschetz dual to $h_{i}\gamma_{i}$, and the curves $\{\gamma_{i}\}$ are all disjoint, one deduces that $p(\mathrm{Twist}(\mathcal{N},\omega))$ is isotropic. We record the following consequence for use later; see also [Wri15, Section 8].
Proposition 5.4.
Let $\omega\in\mathcal{N}$ be horizontally periodic. Then
$\dim_{\mathbb{R}}p(\mathrm{Twist}(\omega,\mathcal{N}))\leq\textrm{rank}(\mathcal{N}).$
Combining Lemma 5.2 and Proposition 5.3 allows us to easily identify orbit closures of twist tori associated to pants decompositions.
Corollary 5.5.
Let $\mathcal{Q}$ be a component of a stratum of holomorphic quadratic differentials on a closed genus $g$ surface and let $\mathcal{L}$ be an AIS. Suppose that $\mathcal{L}$ contains a surface $q$ with 3g33g-3 horizontal cylinders and that all cylinder deformations of these horizontal cylinders remain in $\mathcal{L}$. Then $\mathcal{L}=\mathcal{Q}$.
Proof.
No holomorphic quadratic differential on a genus $g$ surface can have more than 3g33g-3 horizontal cylinders, as their core curves must all be disjoint and homotopically distinct. Thus $q$ contains as many horizontal cylinders as possible for any surface in $\mathcal{Q}$ (hence for any surface in $\mathcal{L}$). By hypothesis, all of these cylinders can be twisted individually to stay in $\mathcal{L}$.
The loci $\widehat{\mathcal{L}}\subset\widehat{\mathcal{Q}}$ of holonomy double covers of surfaces in $\mathcal{L}$ and $\mathcal{Q}$, respectively, are AISs in a stratum of abelian differentials. Every flat cylinder on $q$ lifts to two cylinders on $\hat{q}$, [^16] This can be seen because the holonomy of a curve is the same as its winding number mod 2 with respect to the horizontal line field. hence the holonomy double cover $\hat{q}$ of $q$ has the most horizontal cylinders among all surfaces in $\widehat{\mathcal{Q}}$. The cylinder deformations tangent to $\mathcal{L}$ at $q$ lift to (sums of) cylinder deformations tangent to $\widehat{\mathcal{L}}$ at $\hat{q}$, so
$\dim\mathrm{Twist}(\hat{q},\widehat{\mathcal{L}})=3g-3$
(dimension is preserved because the kernels of transfer maps on (co)homology are always torsion). Using Proposition 5.3 twice, once for $\widehat{\mathcal{L}}$ and once for $\widehat{\mathcal{Q}}$,
[…]
By Lemma 5.2, we conclude that $\widehat{\mathcal{L}}=\widehat{\mathcal{Q}}$, hence that $\mathcal{L}=\mathcal{Q}$, establishing the result. ∎
5.2. Proof of the main theorem
At this point, we can assemble and prove a more detailed version of our main Theorem 1.2. For the reader’s convenience, we quickly recall the setup and the statement below.
[…]
Recall our notation that if $\mathcal{Q}$ is a component of a stratum of quadratic differentials, then $\nu_{\mathcal{Q}}^{1}$ is the Masur–Smillie–Veech probability measure on its unit area locus $\mathcal{Q}^{1}$2.1).
Theorem 5.6.
Fix any pants decomposition $\gamma$, any $\bm{\ell}\in\mathbb{R}^{3g-3}_{>0}$, and any $\mathbf{h}\in\mathbb{R}^{3g-3}_{>0}$ such that $\bm{\ell}\cdot\mathbf{h}=1$. There is a set $Z\subset\mathbb{R}$ of zero density such that
$\lim_{\begin{subarray}{c}t\to\infty\\ t\notin Z\end{subarray}}\mu_{\gamma}(e^{t}\bm{\ell},e^{-t}\mathbf{h})=\mathcal{O}^{*}\nu_{\mathcal{Q}}^{1},$
where $\mathcal{Q}$ is the (connected) stratum $\mathcal{QM}_{g}(1^{4g-4-2\nabla_{\gamma}(\bm{\ell})},2^{\nabla_{\gamma}(\bm{\ell})};\varepsilon)$, and $\varepsilon=1$ if and only if $\nabla_{\gamma}(\bm{\ell})=2g-2$ and $(\gamma,\mathsf{Y})$ is jointly orientable.
Proof of Theorem 1.2 given Theorem 5.6.
We only need to justify the two numbered items. Mirakhani proved that $\mathcal{O}^{*}\nu_{\mathcal{Q}}^{1}=\mu_{\operatorname{Mirz}}/b_{g}$ when $\mathcal{Q}$ is the principal stratum of quadratic differentials [Mir08, Theorems 1.3 and 1.4]. This gives item (1) of Theorem 1.2.
For item (2), note that for any stratum component $\mathcal{Q}$, the measure $\nu_{\mathcal{Q}}^{1}$ is ergodic for the $P$-action. Any two distinct components of strata yield distinct Masur–Smillie–Veech measures. By ergodicity and Theorem 2.3, if $\mathcal{Q}$ is not the principal stratum, then $\mathcal{O}^{*}\nu_{\mathcal{Q}}^{1}$ and $\mu_{\operatorname{Mirz}}$ are mutually singular. ∎
Proof of Theorem 5.6.
[…]
Claim 5.7.
With all notation as above, we have
$\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})}=\mathcal{Q}^{1}\mathcal{M}_{g}(1^{4g-4-2\nabla_{\gamma}(\bm{\ell})},2^{\nabla_{\gamma}(\bm{\ell})};\varepsilon)$
where $\varepsilon=+1$ if $(\gamma,\mathsf{Y})$ is jointly orientable, and $\varepsilon=-1$ if not.
Proof.
The spine $\mathsf{Y}$ has 2g22g-2 components, $\nabla_{\gamma}(\bm{\ell})$ of which have a unique 4-valent vertex and the rest of which have two trivalent vertices. Thus, we can identify the ambient stratum component containing the twist torus using two facts: that $k$-pronged singularities of $\mathcal{O}_{\gamma}(X)$ correspond to zeros of order k2k-2 of $\mathcal{O}(X,\mathbf{h}\cdot\gamma)$ (Theorem 2.3) and that joint orientability of $(\gamma,\mathsf{Y})$ is equivalent to $\mathcal{O}(X,\mathbf{h}\cdot\gamma)$ being the square of an abelian differential (Lemma 4.2). Connectivity of these strata is a consequence of the classification of connected components of strata [KZ03, Lan08].
The argument presented in the proof of Proposition 4.5 gives that $\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})}=\overline{Pq_{0}}$ for generic $q_{0}\in\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$. By Theorem 4.4, $\overline{Pq_{0}}$ is the unit length locus of an AIS $\mathcal{L}$, and since $\overline{Pq_{0}}\supset\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$, Corollary 5.5 implies $\mathcal{L}$ must be the entire ambient stratum. ∎
By Corollary 4.9 and (5), we have that these flat twist tori equidistribute to the Masur–Smillie–Veech probability measure $\nu_{\mathcal{Q}}^{1}$ on the ambient stratum outside a set $Z$ of zero density:
[…]
Applying Theorem 1.3, we therefore get that
[…]
This is what we wanted to show. ∎
6. Other twist tori
One of the pivotal steps in our proof of Theorem 1.2 relied on the fact that no flat surface can have more than 3g33g-3 horizontal cylinders (Corollary 5.5), which makes it seem like pants decompositions are special. In this section, we demonstrate that this phenomenon is actually much more general. Using more tools from Teichmüller dynamics, we show the equidistribution of more families of flat twist tori to their ambient strata (and to other AISs), which implies the equidistribution of other hyperbolic twist tori via Theorem 1.3.
Apart from the identification of orbit closures, the proofs of the results in this section are identical to that of Theorem 1.2 and are therefore omitted. The examples in this section are meant to be illustrative, not exhaustive — we invite the reader to import their favorite techniques to this setting.
6.1. Regular plumbing fixtures
As a motivating example, let us consider the following family of generalizations of our problem. We emphasize, however, that all of the computations in this section are much more general and apply to pushforwards by the dilation flow of any of the tori defined in Section 3.
For $p\geq 3$ and $L>0$, we define the regular plumbing fixture with $p$ holes to be the unique hyperbolic metric on a $p$-holed sphere with totally geodesic boundaries of length $L$ and an (orientation-preserving) dihedral symmetry group of order $2p$. Equivalently, this is the surface corresponding to the ribbon graph obtained by gluing together $p$ intervals of length $L/2$ at their vertices. see Figure 5.
Figure 5. A regular plumbing fixture on a $7$-holed sphere.
Suppose that $\gamma$ is a multicurve on $S$ such that each component $\Sigma_{1},\ldots,\Sigma_{k}$ of $S\setminus\gamma$ has genus 0; let $p_{i}$ denote the number of boundary components of $\Sigma_{i}$. Taking each component to be a regular plumbing fixture with boundary lengths $L$, we get a twist torus $\mathbb{T}_{\gamma}(L)\subset\mathcal{M}_{g}$ whose corresponding metric ribbon graph spine $\mathsf{X}=\mathsf{X}(p_{1},\ldots,p_{k})$ has $k$ components, the $i^{\text{th}}$ of which has two vertices of valence $p_{i}$ and edges of length $L/2$. Paralleling Conjecture 1.1, one can ask: what is the limiting distribution of these tori in $\mathcal{M}_{g}$ as $L\to\infty$?
For any choice of heights $\mathbf{h}$, the corresponding flat twist torus $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{X},\mathbf{h})$ parametrizes surfaces with cone points of angle $p_{1}\pi,p_{1}\pi,p_{2}\pi,p_{2}\pi,\ldots,p_{k}\pi,p_{k}\pi$, so unless all $p_{i}=3$, the limiting distribution of the lifted hyperbolic twist tori $\mathcal{P}\mathbb{T}_{\gamma}(\mathsf{X},\mathbf{h})\subset\mathcal{P}^{1}\mathcal{M}_{g}$ will not be Mirzakhani measure. Moreover, since the orbit closure of $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{X},\mathbf{h})$ is contained in a non-principal stratum, the possible limiting distributions on $\mathcal{M}_{g}$ may be qualitatively different from Mirzakhani measure. For example, the support of the measure may not be (coarsely) dense in $\mathcal{P}^{1}\mathcal{M}_{g}$ (compare [CF24b, Remark 1.3]).
Whenever any $p_{i}$ is even, one can check that the differentials $q\in\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{X},\mathbf{h})$ do not have the maximal number of cylinders for their ambient stratum. For example, if all $p_{i}=4$ then $\gamma=2g-2$, but we have seen in the previous section that there are differentials in the stratum $\mathcal{Q}\mathcal{M}_{g}(2^{2g-2};\varepsilon)$ with an entire pants decomposition worth of horizontal cylinders. Thus we will need to employ different techniques.
6.2. Jointly orientable constructions
We first focus on the case when $(\gamma,\Gamma)$ is jointly orientable, that is, when the twist torus lives inside a stratum of (squares of) abelian differentials. Building on work of Mirzkakhani and Wright on “full rank” AISs [MW18], recent deep work of Apisa and Wright classifies the “large” AISs in these strata.
Following [AW23], say that an AIS $\mathcal{L}$ in a stratum of genus $g$ abelian differentials has high rank if
$\textrm{rank}(\mathcal{L})\geq\frac{g}{2}+1.$